one card is selected at random from a deck of cards. Determine the probability that the card selected is a 5 or a 7
Answer by edjones(8007)

In a 52 card deck there are 4 fives and 4 sevens.
4+4=8
8/52=2/13 the probability that the card selected is a 5 or a 7.
.
Ed